OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 4 to 9. If the perimeter of the larger rectangle is 54 inches, what is the perimeter of the smaller rectangle? (Points : 1) 121.5 inches 36 inches 24 inches 13.5 inches

OpenStudy (johnweldon1993):

Well you know the ratio is \[\large \frac{4}{9} \] and they give you a perimeter for the larger rectangle...so you can do a comparison... \[\large \frac{4}{9} = \frac{x}{54}\] and just solve for 'x'

OpenStudy (anonymous):

ok thanks